# Investigating Twin Star Equation of States in Light of Recent Astrophysical Observations

### v1

@@ I Introduction
The pursuit of understanding the equation of state (EoS) of matter inside a Neutron Star (NS) is a traditional one [1, 2, 3]. It remains an important revisit as we continue to observe more of the universe with increasingly advanced technologies, particularly by measuring these compact objects and witnessing the violent events involving them. NSs possess strong gravity such that the matter at the core is compressed beyond nuclear saturation density $\rho_{\mathrm{sat}}\simeq 2.7\times 10^{14}~\mathrm{g/cm}^{3}$ [4]. The unavailability of such circumstances on earth-based environments makes NSs the natural laboratories in the universe to study the behaviour of matter at low temperatures and high densities [5, 6, 7, 8]. At such extreme conditions, a phase transition (PT) from hadronic matter (HM) to quark matter (QM) is expected from Quantum Chromodynamics (QCD) [9].
@@
TSs are the direct manifestation of strong hadron-to-quark PT. Observation of such candidates of compact objects can confirm the occurrence of PT in the astrophysical scenarios and severely constrain the EoS [81, 66, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], especially if detections are made in the high-mass regime [95, 96, 97, 98]. On the other hand, the formation of TSs in realistic scenarios and their dynamical evolution remain a serious pursuit [99, 100]. Several studies have presented the possibility of realistic EoSs that can construct such species [101, 102, 103, 81, 104, 105, 106, 107, 108, 109, 110, 111, 112]. A few works have studied the oscillation modes [113, 114, 115, 116, 117], QCD trace anomaly [118], rotation [119, 120, 121, 122], magnetic fields [123, 124] and temperature effects [125, 126, 127] on these compact objects.
@@
Figure 1: [Left] Diagram indicating the different parts of EoS, that are controlled by different parameters—transition energy density ($e_{\mathrm{tr}}$), transition pressure ($P_{\mathrm{tr}}$), the strength of discontinuity in energy density ($\Delta e$) and constant speed of sound for quark matter EoS ($c^{2}_{\mathrm{s,QM}}$). [Right] Schematic diagram of MMRR sequence of an EoS that constructs TSs. The second (purple) and third (orange) stable branches are separated by a sequence of unstable branch (black dashed curve). The maximum mass on the second and third branches ($M_{\mathrm{TOV,2}}$ and $M_{\mathrm{TOV,3}}$) is indicated by purple and red stars, respectively. The sequence of unstable stars begins beyond these stellar configurations. The minimum mass on the third branch is indicated by a red circle. The twin region is defined as the mass range in which equal-mass stellar models are present on both stable branches.
@@ II.1 Equation of State
  • Category II: $M_{\mathrm{TOV,2}}\geq 2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\leq 2.0~M_{\odot}$, particularly associated with larger $\Delta e$.
  • Category III: $M_{\mathrm{TOV,2}}\leq 2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\geq 2.0~M_{\odot}$.
  • Category IV: $M_{\mathrm{TOV,2}}\leq 1.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\leq 2.0~M_{\odot}$, particularly associated with low $P_{\mathrm{tr}}$.
@@ Acknowledgements
It is a pleasure to thank Sagnik Chatterjee for his crucial insights into this work. The authors also express their gratitude to the Deucalion HPC in Portugal for its support in the Advanced Computing Project 2025.00067.CPCA A3, RNCA (Rede Nacional de Computação Avançada), financed by the FCT (Fundação para a Ciência e a Tecnologia, IP, Portugal). The test setups were completed in the National Supercomputing Mission (NSM) resource ‘PARAM Ganga’ at IIT Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India. The small-scale tests were conducted at the Bhaskara HPC in IISER Bhopal. SH and RM acknowledge the SERB grant: CRG/2022/000663. This research work is done using the packages—Numpy [144], SciPy [145], Matplotlib [146], Seaborn [147], Jupyter [148], Numba [149] and OpenCV [150].

### v2

@@ I Introduction
The pursuit of understanding the equation of state (EoS) of matter inside a Neutron Star (NS) is a traditional one [106, 115, 41]. It remains an important revisit as we continue to observe more of the universe with increasingly advanced technologies, particularly by measuring these compact objects and witnessing the violent events involving them. NSs possess strong gravity such that the matter at the core is compressed beyond nuclear saturation density $\rho_{\mathrm{sat}}\simeq 2.7\times 10^{14}~\mathrm{g/cm}^{3}$ [70]. The unavailability of such circumstances on earth-based environments makes NSs the natural laboratories in the universe to study the behaviour of matter at low temperatures and high densities [114, 17, 35, 25]. At such extreme conditions, a phase transition (PT) from hadronic matter (HM) to quark matter (QM) is expected from Quantum Chromodynamics (QCD) [159]. Violent astrophysical events, such as supernova explosions [155, 176, 102, 97, 109, 61, 62, 175, 92] and binary NS mergers [135, 171, 145, 60, 30, 23, 168, 100, 93, 64, 65, 85, 141, 39, 90], are predicted to reveal such phenomena.
@@
TSs are the direct manifestation of strong hadron-to-quark PT. Observation of such candidates of compact objects can confirm the occurrence of PT in the astrophysical scenarios and severely constrain the EoS [174, 75, 130, 134, 167, 44, 46, 120, 111, 12, 82, 119, 43, 32, 10], especially if detections are made in the high-mass regime [28, 74, 45, 11]. On the other hand, the formation of TSs in realistic scenarios and their dynamical evolution remain a serious pursuit [59, 138, 38]. Several studies have presented the possibility of realistic EoSs that can construct such species [8, 178, 96, 174, 158, 157, 113, 179, 180, 172, 112, 118, 177, 21]. A few works have studied the oscillation modes [144, 73, 149, 154, 181], modified gravity [124, 123], QCD trace anomaly [98], rotation [29, 52, 27, 139], magnetic fields [72, 161], temperature [125, 36, 37] and dark matter effects [105, 49] on these compact objects.
@@
Figure 1: [Left] Diagram indicating the different parts of EoS, that are controlled by different parameters—transition energy density ($e_{\mathrm{tr}}$), transition pressure ($P_{\mathrm{tr}}$), the strength of discontinuity in energy density ($\Delta e$) and constant speed of sound for quark matter EoS ($c^{2}_{\mathrm{s,QM}}$). [Right] Diagram of MMRR sequence of an EoS that constructs TSs. The second (purple) and third (orange) stable branches are separated by a sequence of unstable branch (black dashed curve). The maximum mass on the second and third branches ($M_{\mathrm{TOV,2}}$ and $M_{\mathrm{TOV,3}}$) is indicated by purple and red stars, respectively. The sequence of unstable stars begins beyond these stellar configurations. The minimum mass on the third branch is indicated by a red circle. The twin region is defined as the mass range in which equal-mass stellar models are present on both stable branches.
@@ II.1 Equation of State
  • Category II: $M_{\mathrm{TOV,2}}\geq 2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}<2.0~M_{\odot}$, particularly associated with larger $\Delta e$.
  • Category III: $1.0~M_{\odot}\leq M_{\mathrm{TOV,2}}<2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\geq 2.0~M_{\odot}$.
  • Category IV: $M_{\mathrm{TOV,2}}<1.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\geq 2.0~M_{\odot}$, particularly associated with low $P_{\mathrm{tr}}$.
@@ IV Conclusions
Finally, we would like to emphasise the criticality of the existence of TSs of category II. We note that in a large ensemble of EoSs, only a small fraction of EoSs with TSs of category II are allowed. In ref. [32], this category of TSs is shown to be unlikely based on a Bayesian study. These inferences are based solely on hydrostatic equilibrium solutions and fail to account for the dynamical response of TSs to astrophysical perturbations. In a dedicated study [86], we show that, under small perturbations, TSs of category II either favour migrating to the hadronic stars on neighbouring branch or collapsing into black holes.
@@ Acknowledgements
It is a pleasure to thank Sagnik Chatterjee and Luciano Rezzolla for their crucial insights into this work. The authors also express their gratitude to the Deucalion HPC in Portugal for its support in the Advanced Computing Project 2025.00067.CPCA A3, RNCA (Rede Nacional de Computação Avançada), financed by the FCT (Fundação para a Ciência e a Tecnologia, IP, Portugal). The test setups were completed in the National Supercomputing Mission (NSM) resource ‘PARAM Ganga’ at IIT Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India. The small-scale tests were conducted at the Bhaskara HPC in IISER Bhopal. SH and RM acknowledge the SERB grant: CRG/2022/000663. This research work is done using the packages—Numpy [87], SciPy [169], Matplotlib [94], Seaborn [170], Jupyter [104], Numba [110] and OpenCV [34].